OpenStudy (anonymous):

which expression is equivalent to 4(2x+1)-(-6x) 14x+4 8x-2 2x+4 -14x-4

OpenStudy (anonymous):

start by using the distributive law \[4(2x+1)-(-6x)=4\times 2x+4\times 1+6x\] then go to \[8x+4+6x\] and finally combine like terms

OpenStudy (anonymous):

explain to me how you combine like terms

OpenStudy (anonymous):

8 bottles of beer and 6 bottles of beer is a total of 14 bottles of beer

OpenStudy (anonymous):

similarly \(8x+6x=14x\)

OpenStudy (anonymous):

as long as you understand that \(8x\) means "eight times \(x\)" or \(8\times x\) then combining like terms is a matter of adding up the coefficients of the like terms